# The dark matter content of Milky Way dwarf spheroidal galaxies: Draco, Sextans and Ursa Minor

### v1

@@ 4 Results
In this section, we present our results for both single-population and chemodynamical models. For each model, we show best-fit model parameters as well as best-fit velocity moments compared with the real observations. For the chemodynamical model, we also present the spatial, kinematical and metallicity distributions of the two populations separated by the best model. In the end, we calculate the astrophysical J and D factors for dark matter indirect detections.
@@ 4.1 Best constraints and performance of the single-population model
Object $\rho_{s}[\mathrm{M_{\odot}\,pc^{-3}}]$ $r_{s}[\mathrm{kpc}]$ $\gamma$ $i[\mathrm{deg}]$ $\lambda$ $\kappa$
Draco $0.033^{+0.176}_{-0.028}$ $2.12^{+3.87}_{-1.35}$ $0.82^{+0.36}_{-0.42}$ $77^{+9}_{-10}$ $0.65^{+0.19}_{-0.17}$ $0.43^{+0.21}_{-0.15}$
Sextans $0.016^{+0.010}_{-0.008}$ $6.22^{+2.50}_{-2.64}$ $0.29^{+0.23}_{-0.13}$ $64^{+14}_{-8}$ $0.71^{+0.33}_{-0.22}$ $0.38^{+0.19}_{-0.16}$
UMi $0.023^{+0.022}_{-0.013}$ $4.70^{+3.17}_{-2.32}$ $0.35^{+0.24}_{-0.17}$ $82^{+5}_{-7}$ $1.32^{+0.19}_{-0.15}$ $-0.15^{+0.11}_{-0.12}$
@@
The likelihood contours for different parameters combinations in the single-population model are shown in Figure 6 for Draco only. The trends are similar for Sextans and UMi. Our model imposes weaker constraints on the inclination angle ($i$) compared to other parameters, which is indicated by its wide posterior distribution. The rotation parameter ($\kappa$) shows a weak correlation with the potential parameters ($d_{1}$, $d_{2}$ and $\gamma$), whereas there are strong degeneracies among the potential parameters themselves. The correlation between the velocity anisotropy ($\lambda$) and the other potential parameters is also present, which is the mass-anisotropy degeneracy.
The mass-anisotropy degeneracy can be weakened if we have precise proper motion measurements. However, at the distances of the three dSphs used in this paper, their Gaia proper motions used in our constraints are subject to very large errors, and as we have mentioned, including or not including Gaia proper motions in our analysis leads to very similar constraints. More precise proper motions can possibly be achieved by combining Gaia and previous observations by the Hubble Space Telescope (e.g. del Pino et al., 2022), but this is only available for a small number of stars in more central regions, and thus in this study we focus on using only Gaia proper motions.
@@
In Figure 7, the observed and best-fit first (mean velocity) and second (velocity dispersion) velocity moments are shown for the three dSphs. The mean velocity and velocity dispersion profiles are calculated from the LOSVs along $z^{\prime}$. In general, the velocity moments predicted by the best models are consistent with data, with the velocity dispersion profiles show good matches in amplitudes. The model can also recover the general shapes in the velocity dispersion profiles. In the velocity dispersion profile along the major axis of Draco and Sextans (second panels from the left in the top and middle rows), however, a few green dots drop below the model curves at $R^{\prime}\sim 10-20\ \mathrm{arcmin}$. This is likely driven by the outer most green dots, that causes the model to match the velocity dispersion in the very outer region, sacrificing the match at $R^{\prime}\sim 10-20\ \mathrm{arcmin}$.
@@ 4.2 Best constraints and performance of the chemodynamical model
The full list of best-fit parameters for the chemodynamical model and for the three dSphs is provided in Table 5. Again we note large uncertainties in the inclination angle, $i$. Moreover, for the fractions $h_{1}$ and $h_{2}$ of metal-poor stars in red and blue templates (see Equation 13), they are close to 0 and 1, respectively. This shows that our template maps, though created through simpler hard cuts in stellar metallicities, lead to similar divisions in the stellar populations as the soft divisions according to posterior distributions of the best constrained model, which takes into account the joint spatial, kinematical and metallicity probabilities.
We also present the best constrained model parameters for the chemodynamical model by green symbols in Figure 5. The best-fit model parameters for the single-population and chemodynamical models are consistent with each other within the errors. For the tracer parameters ($\lambda$ and $\kappa$), constraints of two populations in the chemodynamical model feature larger errorbars than those of the single-population model. This is because the sample size of each population is smaller than that of the total sample.
Object $\rho_{s}[\mathrm{M_{\odot}pc^{-3}}]$ $r_{s}[\mathrm{kpc}]$ $\gamma$ $i[\mathrm{deg}]$ $\lambda^{\mathrm{mr}}$ $\lambda^{\mathrm{mp}}$ $\kappa^{\mathrm{mr}}$ $\kappa^{\mathrm{mp}}$
$Z_{0}^{\mathrm{mr}}$ $Z_{0}^{\mathrm{mp}}$ $\sigma_{Z}^{\mathrm{mr}}$ $\sigma_{Z}^{\mathrm{mp}}$ $h_{1}$ $h_{2}$
Draco $0.036^{+0.147}_{-0.028}$ $2.58^{+3.52}_{-1.67}$ $0.71^{+0.34}_{-0.35}$ $79^{+8}_{-9}$ $0.50^{+0.17}_{-0.23}$ $0.99^{+0.29}_{-0.24}$ $0.37^{+0.25}_{-0.24}$ $0.33^{+0.27}_{-0.21}$
$-1.75^{+0.04}_{-0.04}$ $-2.28^{+0.11}_{-0.09}$ $0.24^{+0.03}_{-0.03}$ $0.36^{+0.04}_{-0.04}$ $0.07^{+0.16}_{-0.05}$ $0.91^{+0.07}_{-0.15}$
Sextans $0.019^{+0.011}_{-0.009}$ $5.82^{+2.71}_{-2.39}$ $0.26^{+0.22}_{-0.12}$ $62^{+12}_{-8}$ $1.01^{+0.57}_{-0.32}$ $0.59^{+0.38}_{-0.24}$ $0.10^{+0.41}_{-0.46}$ $0.41^{+0.25}_{-0.23}$
$-1.71^{+0.05}_{-0.05}$ $-2.24^{+0.05}_{-0.06}$ $0.27^{+0.04}_{-0.04}$ $0.50^{+0.03}_{-0.04}$ $0.07^{+0.09}_{-0.05}$ $0.95^{+0.04}_{-0.09}$
UMi $0.023^{+0.015}_{-0.011}$ $5.62^{+2.89}_{-2.53}$ $0.33^{+0.20}_{-0.16}$ $83^{+5}_{-7}$ $1.42^{+0.22}_{-0.23}$ $1.23^{+0.16}_{-0.13}$ $-0.09^{+0.23}_{-0.21}$ $-0.21^{+0.25}_{-0.25}$
$-1.96^{+0.02}_{-0.02}$ $-2.32^{+0.04}_{-0.04}$ $0.26^{+0.02}_{-0.02}$ $0.50^{+0.03}_{-0.02}$ $0.03^{+0.05}_{-0.02}$ $0.89^{+0.08}_{-0.13}$
@@
We show the best constrained dark matter density profiles through both single-population (black) and chemodynamical (orange) models in Figure 8. Given the consistency in the best constrained model parameters in Figure 5, the best-fit dark matter density profiles by the two models are also fully consistent with each other. Dark matter density profiles are best constrained around the half-number radii of tracer populations (dashed vertical lines), in that the associated statistical errors indicated by the shaded regions are smallest. This agrees with the consensus that dynamical modeling generally imposes strongest constraints on the total mass enclosed in the half-number radius of tracers (e.g. Wolf et al., 2010; Walker & Peñarrubia, 2011; Wang et al., 2015; Han et al., 2016a, b; Genina et al., 2020; Wang et al., 2022; Li et al., 2022; Shi et al., 2024).
According to Figure 8, it can be seen that the half-number radii of the metal-rich and metal-poor populations sit on either sides of the half-number radii of the total population, with the metal-rich population having smaller half-number radii, but the difference is not very large. In principle, chemodynamical modeling with two distinct populations is expected to have tighter constraints at two different half-number radii of the two underlying populations, hence leads to better constraints, but Figure 8 shows that the best constraints from single-population and chemodynamical models in our analysis are very similar. This is perhaps due to the small difference in the half-number radii of the metal-rich and metal-poor populations.
Among the three dSphs, Draco has the steepest inner density profile. It favors an inner dark matter density profile $\gamma=1$ within its $1\sigma$ confidence intervals, which is consistent with cusped NFW profile. On the other hand, dark matter halos of Sextans and UMi have lower central densities and their inner profiles are closer to cores. In our samples, the inner density slope $\gamma$ of the dark matter halo features a diversity rather than a universal value. We will back discuss this point further in Section 5 below.
@@
The two populations also differ in their global velocity distributions. Figure 10 more clearly shows the first (mean LOSV) and second (LOSV dispersion) velocity moments of the two populations, in a similar manner as Figure 7 for the single-population model, but the two stellar populations are shown in the two left (metal-poor) and two right (metal-rich) columns separately. Here we only show velocity moment profiles along major axes.
@@
Figure 9: The spatial, metallicity and velocity distributions for the metal-rich (red points/histograms) and metal-poor (blue points/histograms) populations in three dSphs recovered by the chemodynamical model. Each sub plot has five panels. Top histogram: the spatial distributions of two populations. Middle-left scatter panel: $[\mathrm{Fe}/\mathrm{H}]$ versus the projected semimajor radius $x^{\prime}$. Middle-right histogram: metallicity distributions of two populations. Bottom-left scatter panel: the LOSV ($v_{z}^{\prime}$) of each star, versus the projected semimajor radius $x^{\prime}$. Bottom-right histogram: the global velocity distributions of two populations. The division of the two populations are based on the posterior probability of our best-fit model (see Equation 17), with metal-rich and metal-poor populations have this probability greater than 0.5 ($P_{i}^{\prime\,\mathrm{metal-rich}}>0.5$). In the two right panels showing the histogram distributions, histograms are based on the data, while solid smooth curves with the same colors are model predictions. Grey dashed horizontal lines in middle panels denote the hard cuts in metallicity to create the red and blue spatial templates.
@@
For Draco, the two populations identified by the best-fit models (symbols with error bars) show similar mean velocity profiles, which indicate rotations around the minor axes but with different amplitudes. The best-fit model predicts different signs of rotations for the two populations in Sextans, which was also seen in Sculptor (Zhu et al., 2016a). For UMi, rotation is only seen in the metal-poor population. However, the two populations show distinct velocity dispersion profiles in all three dSphs. The metal-poor populations are usually kinematically hotter, with negative gradients in their dispersion profiles that decrease with radius. On the other hand, velocity dispersion profiles of the metal-rich populations are kinematically colder and nearly flat, only slightly declining in outer regions for Sextans and UMi. These distinct properties of velocity dispersions of two populations are also consistent with those reported in previous studies (e.g. Pace et al., 2020), in that the younger metal-rich population may involve more later formed stars in central regions which have more coherent motions, hence the velocity dispersions are decreased. On the other hand, the older metal-poor population is more relaxed.
@@
In general, the predicted mean velocity profiles by the best constrained model agree well with the data. Unfortunately, the velocity dispersion profiles are poorly fit for the metal-poor population of Draco, both populations of Sextans and the metal-poor population of UMi. These poor fittings are mainly caused by the relatively large difference between the velocity dispersions of the two populations. For all three dSphs, the central velocity dispersions of the metal-poor populations are nearly a factor of two larger than those of the metal-rich populations. If different populations have different velocity anisotropy parameters, $\lambda$, it may cause some differences in the velocity dispersion for the LOSV component. But as we have tested different values of $\lambda$ to predict the velocity dispersions using Jeans equations, the difference is difficult to be as large as a factor of two. Moreover, the significantly higher velocity dispersion in metal-poor populations cannot be due to the contamination by RR Lyrae stars either, as we have excluded the entire horizontal branch (see Section 2). Moreover, we have tried to exclude a few stars with large LOSV variations based on their multi-epoch observations, which are more likely binaries, but the velocity dispersions almost do not change. We have also checked that the different LOSV dispersions of the two populations cannot be due to their different LOSV errors either.
@@
The nearly factor of two difference in the velocity dispersions of the metal-rich and metal-poor populations may reflect that the dSph system is deviating from the steady-state assumption behind the Jeans equations. On the other hand, for multiple tracer populations bound together and moving in the same potential, if all tracers are in steady state, we do not expect such prominent differences in their velocity dispersions. Thus the solutions to Jeans equations under the steady-state tracer assumptions are impossible to reproduce such a prominent difference between different stellar populations moving in the same potential. Hence impossible to simultaneously achieve ideal fits in the observed velocity dispersion profiles of both populations.
@@
In particular, for Draco and UMi, the metal-rich populations in the rightmost panels dominate in sample size (see Figure 9 above), and thus the models tend to match better the metal-rich population. For Sextans, the metal-rich and metal-poor populations are more comparable in sample size, and the best-fit model tends to reconcile the large difference in velocity dispersion profiles of two populations, resulting in over-estimated dispersions for the metal-rich population, and under-estimated dispersions for the metal-poor population.

### v2

@@ 4 Results
In this section, we present our results for both single-population and chemodynamical models. For each model, we show best-fit model parameters as well as best-fit velocity moments compared with the real observations. For the chemodynamical model, we also present the spatial, kinematical and metallicity distributions of the two populations separated by the best model. In the end, we calculate the astrophysical J and D factors for indirect dark matter detections.
@@ 4.1 Best constraints and performance of the single-population model
Parameter Draco Sextans UMi
$\rho_{s}\ [\mathrm{M_{\odot}\,pc^{-3}}]$ $033^{+0.176}_{-0.028} 0.0$$16^{+0.010}_{-0.008} 0.0$$023^{+0.022}_{-0.013}$
$r_{s}\ [\mathrm{kpc}]$ $2.12^{+3.87}_{-1.35}$ $6.22^{+2.50}_{-2.64}$ $4.70^{+3.17}_{-2.32}$
$\gamma$ $0.82^{+0.36}_{-0.42}$ $0.29^{+0.23}_{-0.13}$ $0.35^{+0.24}_{-0.17}$
i[deg]i\ $\mathrm${deg}] $77^{+9}_{-10}$ $64^{+14}_{-8}$ $82^{+5}_{-7}$
$\lambda$ $0.65^{+0.19}_{-0.17}$ $0.71^{+0.33}_{-0.22}$ $1.32^{+0.19}_{-0.15}$
$\kappa$ $0.43^{+0.21}_{-0.15}$ $0.38^{+0.19}_{-0.16}$ $0.15^{+0.11}_{-0.12}$
@@
The likelihood contours for different parameter combinations in the single-population model are shown in Figure 6 for Draco only. The trends are similar for Sextans and UMi. Our model imposes weaker constraints on the inclination angle ($i$) compared to other parameters, which is indicated by its wide posterior distribution. The rotation parameter ($\kappa$) shows a weak correlation with the potential parameters ($d_{1}$, $d_{2}$ and $\gamma$), whereas there are strong degeneracies among the potential parameters themselves. The correlation between the velocity anisotropy ($\lambda$) and the other potential parameters is also present, which is the mass-anisotropy degeneracy.
The mass-anisotropy degeneracy can be weakened if we have precise proper motion measurements. However, at the distances of the three dSphs used in this paper, their Gaia proper motions used in our constraints are subject to very large errors, and as we have mentioned, including or not including Gaia proper motions in our analysis leads to very similar constraints. More precise proper motions might be achieved by combining Gaia and previous observations by the Hubble Space Telescope (HST, e.g. del Pino et al., 2022), but this is only available for a small number of stars in the center. Thus in this study we only use Gaia proper motions. In fact, Vitral et al. (2024) recently combined precise proper motion measurements for about 400 stars in Draco from HST observations with LOSV measurements from Walker et al. (2023), and performed modeling similar to jam. Their inferred inner dark matter density slope is fully consistent with our result.
@@
The observed and best-fit first (mean velocity) and second (velocity dispersion) velocity moments are shown in Figure 7. The mean velocity and velocity dispersion profiles are calculated from the LOSVs along $z^{\prime}$. In general, the velocity moments predicted by the best models are consistent with data, with the velocity dispersion profiles show good matches in amplitudes. The model can also recover the general shapes in the velocity dispersion profiles. In the velocity dispersion profile along the major axis of Draco and Sextans (second panels from the left in the top and middle rows), however, a few green dots drop below the model curves at $R^{\prime}\sim 10-20\ \mathrm{arcmin}$. This is likely driven by the outer most green dots, that causes the model to match the velocity dispersion in the very outer region, sacrificing the match at $R^{\prime}\sim 10-20\ \mathrm{arcmin}$.
@@ 4.2 Best constraints and performance of the chemodynamical model
The full list of best-fit parameters for the chemodynamical model is shown in Table 3. For the fractions $h_{1}$ and $h_{2}$ of metal-poor stars in red and blue templates (see Equation 13), they are close to 0 and 1, respectively. This shows that our template maps, though created through simpler hard cuts in stellar metallicities, lead to similar divisions in the stellar populations as the soft divisions according to posterior distributions of the best constrained model, which takes into account the joint spatial, kinematical and metallicity probabilities.
We also present the best constrained parameters for the chemodynamical model by green symbols in Figure 5. The best-fit parameters for the single-population and chemodynamical models are consistent with each other. For the tracer parameters ($\lambda$ and $\kappa$), constraints of two populations in the chemodynamical model feature larger errorbars than those of the single-population model. This is because the sample size of each population is smaller than that of the total sample.
Parameter Draco Sextans UMi
$\rho_{s}\ [\mathrm{M_{\odot}pc^{-3}}]$ $036^{+0.147}_{-0.028} 0.0$$19^{+0.011}_{-0.009} 0.0$$023^{+0.015}_{-0.011}$
$r_{s}\ [\mathrm{kpc}]$ $2.58^{+3.52}_{-1.67}$ $5.82^{+2.71}_{-2.39}$ $5.62^{+2.89}_{-2.53}$
$\gamma$ $0.71^{+0.34}_{-0.35}$ $0.26^{+0.22}_{-0.12}$ $0.33^{+0.20}_{-0.16}$
i[deg]i\ $\mathrm${deg}] $79^{+8}_{-9}$ $62^{+12}_{-8}$ $83^{+5}_{-7}$
$\lambda^{\mathrm{mr}}$ $0.50^{+0.17}_{-0.23}$ $1.01^{+0.57}_{-0.32}$ $1.42^{+0.22}_{-0.23}$
$\lambda^{\mathrm{mp}}$ $0.99^{+0.29}_{-0.24}$ $0.59^{+0.38}_{-0.24}$ $1.23^{+0.16}_{-0.13}$
$\kappa^{\mathrm{mr}}$ $0.37^{+0.25}_{-0.24}$ $10^{+0.41}_{-0.46} 0.0$$09^{+0.23}_{-0.21}$
$\kappa^{\mathrm{mp}}$ $0.33^{+0.27}_{-0.21}$ $0.41^{+0.25}_{-0.23}$ $0.21^{+0.25}_{-0.25}$
$Z_{0}^{\mathrm{mr}}$ $1.75^{+0.04}_{-0.04}$ $1.71^{+0.05}_{-0.05}$ $1.96^{+0.02}_{-0.02}$
$Z_{0}^{\mathrm{mp}}$ $28^{+0.11}_{-0.09} 2.2$$24^{+0.05}_{-0.06}$ $2.32^{+0.04}_{-0.04}$
$\sigma_{Z}^{\mathrm{mr}}$ $0.24^{+0.03}_{-0.03}$ $0.27^{+0.04}_{-0.04}$ $0.26^{+0.02}_{-0.02}$
$\sigma_{Z}^{\mathrm{mp}}$ $0.36^{+0.04}_{-0.04}$ $0.50^{+0.03}_{-0.04}$ $0.50^{+0.03}_{-0.02}$
$h_{1}$ $07^{+0.16}_{-0.05} 0.0$$7^{+0.09}_{-0.05} 0.0$$03^{+0.05}_{-0.02}$
$h_{2}$ $0.91^{+0.07}_{-0.15}$ $0.95^{+0.04}_{-0.09}$ $0.89^{+0.08}_{-0.13}$
@@
We show the best constrained dark matter density profiles through both single-population (black) and chemodynamical (orange) models in Figure 8. Given the consistency in the best constrained model parameters in Figure 5, the best-fit dark matter density profiles by the two models are also fully consistent with each other. Dark matter density profiles are best constrained around the half-number radii of tracer populations (dashed vertical lines), where the associated statistical errors indicated by the shaded regions around density profiles are the smallest. This agrees with the consensus that dynamical modeling generally imposes strongest constraints on the total mass enclosed in the half-number radius of tracers (e.g. Wolf et al., 2010; Walker & Peñarrubia, 2011; Wang et al., 2015; Han et al., 2016a, b; Genina et al., 2020; Wang et al., 2022; Li et al., 2022; Shi et al., 2024).
According to Figure 8, it can be seen that the half-number radii of the metal-rich and metal-poor populations sit on either sides of the half-number radii of the total population, with the metal-rich population having smaller half-number radii. In principle, chemodynamical modeling with two distinct populations is expected to have tighter constraints at two different half-number radii of the two underlying populations, hence leads to better constraints, but Figure 8 shows that the best constraints from single-population and chemodynamical models in our analysis are consistent within the errors. This could be limited by the number of available tracer stars used and the relatively large statistical errors of the density profiles, but at least, the agreement between the best constraints of our single-population and two-population chemodynamical models show that our results are self-consistent.
Among the three dSphs, Draco has the steepest inner density profile. It favors an inner dark matter density profile $\gamma=1$ within its $1\sigma$ confidence intervals, which is consistent with cusped NFW profile. On the other hand, dark matter halos of Sextans and UMi have lower central densities and their inner profiles are closer to cores. Across the three dSphs studied in this paper, the inner density slope $\gamma$ of the dark matter halo features a diversity rather than a universal value. We will back discuss this point further in Section 5 below.
@@
The two populations also differ in their global velocity distributions. Figure 10 more clearly shows the first (mean LOSV) and second (LOSV dispersion) velocity moments of the two populations, similarly to Figure 7 for the single-population model, but the two populations are shown in the two left (metal-poor) and two right (metal-rich) columns separately. Here we only show velocity moment profiles along major axes.
@@
Figure 9: The spatial, metallicity and velocity distributions for the metal-rich (red) and metal-poor (blue) populations recovered by the chemodynamical model. Each sub plot has five panels. Top histogram: the spatial distributions of two populations. Middle-left scatter panel: $\mathrm{Fe}/\mathrm{H}]$ versus the projected radius $R^{\prime}$. Middle-right histogram: metallicity distributions of two populations. Bottom-left scatter panel: the LOSV ($v_{z}^{\prime}$) of each star, versus the projected radius $R^{\prime}$. Bottom-right histogram: the global velocity distributions of two populations. The division of the two populations are based on Equation 17, with metal-rich and metal-poor populations having this probability greater than 0.5. The histogram distributions in the two right panels are for the data, while solid smooth curves with the same colors are model predictions. Grey dashed horizontal lines denote the hard cuts in metallicity to create the red and blue spatial templates.
@@
For Draco, the two populations identified by the best-fit models (symbols with error bars) show similar mean velocity profiles, which indicate rotations around the minor axes but with different amplitudes. The best-fit model predicts different signs of rotations for the two populations in Sextans, which was also seen in Sculptor (Zhu et al., 2016a). For UMi, rotation is only seen in the metal-poor population. However, the two populations show distinct velocity dispersion profiles in all three dSphs. The metal-poor populations are usually kinematically hotter, with negative gradients in their dispersion profiles that decrease with radius. On the other hand, velocity dispersion profiles of the metal-rich populations are kinematically colder and more flattened. The distinct velocity dispersions of the two populations are also consistent with previous studies (e.g. Pace et al., 2020), in that the later formed metal-rich population may have more coherent motions, hence the velocity dispersions are decreased. The older metal-poor population is more relaxed.
@@
In general, the predicted mean velocity profiles by the best constrained model agree well with the data. Unfortunately, the velocity dispersion profiles are poorly fit for the metal-poor populations of both Draco and UMi, and for both populations of Sextans. The same is seen in the metal-poor population of Sculptor in Zhu et al. (2016a). These poor fits are mainly caused by the large difference between the velocity dispersions of the two populations. For all three dSphs, the central velocity dispersions of the metal-poor populations are nearly a factor of two larger than those of the metal-rich ones. If the velocity anisotropy parameters, $\lambda$, for different populations differ, it may cause some differences in the velocity dispersions. But we have tested different $\lambda$ to predict the velocity dispersions using Jeans equations, it is difficult to reproduce a difference as large as a factor of two. The significantly higher velocity dispersion in metal-poor populations cannot be due to the contamination by RR Lyrae stars, as we have excluded the entire horizontal branch (see Section 2). Moreover, the effect of binaries cannot account for it either. Among the three dSphs, UMi has multi-epoch observations by DESI for more than six hundred member stars, with a time baseline spanning one year. We have excluded stars with large LOSV variations, defined as the absolute LOSV change between two epoches of observations being greater than a factor 3 of the LOSV error, but the resulting velocity dispersion profiles are very similar to the original ones. We have also checked that the different LOSV dispersions of the two populations cannot be due to their different LOSV errors.
@@
The nearly factor of two difference in the velocity dispersions of the two populations may reflect that the dSph systems deviate from the steady-state assumption behind the Jeans equations, as for multiple tracer populations in the same potential, if all tracers are in steady states, we do not expect such prominent differences in their velocity dispersions. Recent Gaia observations have also indicated that many MW dSphs may be out of equilibrium (Hammer et al., 2023; Yang et al., 2024; Wang et al., 2024). The steady-state Jeans solutions cannot reproduce such a prominent difference between different populations in the same potential. Hence it is difficult to simultaneously achieve ideal fits in the observed velocity dispersion profiles of both populations.
@@
For Draco and UMi, the metal-rich populations in the rightmost panels dominate in sample size (see Figure 9 above), and thus the models tend to match better the metal-rich population. For Sextans, the metal-rich and metal-poor populations are more comparable in sample size, and the best-fit model tends to reconcile the large difference in velocity dispersion profiles of two populations, resulting in over-estimated dispersions for the metal-rich population, and under-estimated dispersions for the metal-poor population.
Deviations from steady-state assumption can introduce biased constraints. Previously, (Wang et al., 2022) applied jam to simulated dwarf galaxies to test the biases. It was found that global expansion/infalling motions can cause deviations from steady states, resulting in over/under-estimated inner densities, but still maintain a good constraint on the total mass within the half-number radii of tracers. In our analysis, we introduce the two-population chemodynamical model because the two distinct half-number radii of the metal-rich and metal-poor populations may help to reduce systematics. Our results are at least self-consistent given the fact that the single-population and chemodynamical models are consistent.